OpenStudy (anonymous):

5. A system of linear inequalities is shown below: x + y ≥ 4 y < 2x – 3 Describe the steps to graph the solution set to the system of inequalities. Be specific. (10 points) @bruhhh @Hero

OpenStudy (anonymous):

@uri @Secret-Ninja @Savannah22 @sarah2001 @Sadworld

OpenStudy (anonymous):

Do you just use intercepts?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

what do you mean by that

OpenStudy (anonymous):

I'm assuming you have to graph this.

OpenStudy (anonymous):

no not for this question

OpenStudy (anonymous):

I'm not really sure about this but one way of doing it is I'd rewrite them as functions. Then I would substitute x with my own domain (1, 2, 3, etc.). Solve to get the value of y. Then proceed to input the coordinates in the chart.

OpenStudy (anonymous):

Another way would be to transpose the first set (x+y > 4) so that x is at one side and the rest are on the other (x>4-y). I'd then replace x in the second equation (y < 2x – 3) => (y < 2(4-y) – 3). I might be wrong.

OpenStudy (anonymous):

CAn you help with another one

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

thanks ok here it is

OpenStudy (anonymous):

4. Part A: Explain why the x-coordinates of the points where the graphs of the equations y=2-x and y = 4x + 3 intersect are the solutions of the equation 2-x = 4x + 3 Hint: Think about graphing (5 points) Part B: Make tables to find the solution to 2-x = 4x + 3. Take the values of x between −3 and 3. (5 points) Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2-x and y = 4x + 3 intersect are the solutions of the equation 2-x = 4x + 3. Hint: Think about graphing (5 points) Part B: Make tables to find the solution to 2-x = 4x + 3. Take the values of x between −3 and 3. Be sure to state the solution. (5 points) y = 2-x y = 4x + 3 x y 1 0 -1 -2 x y 1 0 -1 -2

OpenStudy (anonymous):

Can you ask this in another question? I'd like the extra medal so I can unlock a title :P

OpenStudy (anonymous):

ok